# Feedback Intensity Equalization Algorithm for Multi-Spots Holographic Tweezer

### v2

@@
Feedback Intensity Equalization Algorithm for Multi-Spots Holographic Tweezer
@@
††journal: opticajournal††articletype: Research ArticleThanks to the high degree of adjustability, the holographic tweezer array has been proven to be the best choice for creating an arbitrary geometrical atomic array. In holographic tweezer array experiments, an optical tweezer generated by a spatial light modulator (SLM) usually is used as a static tweezer array. Due to the AC stark effect, the intensity difference of traps will cause different light shifts. So, intensity equalization is very important in a many-body system consisting of single atoms. Here we report a work on studying of intensity equalization algorithm. Through this algorithm, the uniformity of the tweezer can exceed 96 when the array size is bigger than 1000. Our analysis shows that further uniformity requires further optimization of the optical system. The realization of the intensity equalization algorithm is of great significance to quantum computation and simulation experiments based on a single-atom array.
1 Introduction
The holographic tweezer array formed by SLM is a technique that uses interference and diffraction principles to reconstruct the desired optical arrays[1, 2]. There are several protocols to produce the tweezer array, such as, by digital micro-mirror devices (DMD)[3], by Talbot effect[4], by micro lens[5], by meta surface[6]. Compared to the SLM method, the diffraction efficiency of DMD and meta surface are limited, while the arbitrary editing capability is impossible for Talbot and microlens array. The phase pattern of SLM can be set on demand[7, 8], which makes it flexible for producing arbitrary arrays and correcting the aberration[9, 10, 11, 12]. Therefore, the SLM has become the most effective equipment for optical tweezer array system[13, 14]. Optical tweezer arrays have been applied to the cold atom system to trap cold atoms. Due to the collisional-blocked effect, when the tweezer waist is small enough, there will be only one atom at most can be trapped in the tweezer. Combined with rearrangement technology, two - and three-dimensional defect-free atomic arrays have been realized in the last decade[15, 16]. This technique has induced many applications in neutral atom array systems including quantum computation [17]and simulation[18, 19], and quantum metrology[20]. In these experiments, the tweezer arrays made by SLM are used for static tweezer that plays the role of trapping the initial atomic array for sorting and holding the sorted atomic array[21]. According to the calculation formula of the trap depth, the trap depth is proportional to the trap intensity. In the case of different trap depths, the loss probability of the trapped atom is not the same. And different trap depths will lead to different light shifts. So, to prepare an ideal atomic array, an intensity equalized trap array is necessary[21]. Here we report a work of intensity equalization by feedback weighted-Gerchberg-Saxton (GSW) algorithm[22, 23]. Through the imaging of the generated array, we create the corrected target and feed it back to the GSW algorithm. By iterating this feedback process several times, the uniformity of the tweezer array can be improved greatly. The results show that the uniformity can exceed 96% when the array size is over 1000.
2 Principle of Feedback GSW
@@
The core of our algorithm is the GSW algorithm. The algorithm is based on Fourier optics theory and iteratively adjusts the phase distribution on a SLM to achieve the desired intensity distribution on the focal plane. The GSW algorithm optimizes the uniformity and quality of the holographic optical tweezer array by introducing weighted factors. We first generate a desired target. The SLM is placed at the back focal plane (x-plane) of the lens. The reflected wavefront is shaped by a computer-generated hologram (CGH) [24] displayed on the SLM, then forms the tweezer array at u-plane.
@@
According to the imaging principle of Fourier optics, the phase pattern in the x-plane can be calculated as . Due to the phase-only spatial light modulator, here we assign the amplitude to 1 and keep the phase component for inverse fast Fourier-transform(IFFT). A new target graph can be obtained by this IFFT. And we could get the intensity distribution . To approach the desired target, a weighted factor is calculated by the expression below:
@@
After iterating many times, the calculated target will be close to the desired target we input. This algorithm is a standard process of GSW. But in practice, a grating phase needs to be added to shift the target away from the zeroth order. And due to the imperfect optical system, the achieved target is not uniform as the calculated result. To correct this non-uniformity caused by the optical system, we apply another feedback scheme. The basic principle is modifying the input target of GSW according to the non-uniformity target achieved by the GSW algorithm. This so-called feedback GSW algorithm begins with imaging the i-th GSW algorithm and analyzing the uniformity by 2D Gaussian fitting of each spot. The uniformity is calculated by
where and are the maximum and minimum intensity fitting results of i-th array. The (i+1)-th intensity is calculated by
where is the reconstructed target intensity, is the target intensity of last loop. is the average intensity of 2D Gaussian fitting intensities, is the k-th intensity of each spot. G is the weighted factor. Here we apply a constant number as a weighted factor. It results that the best performance is achieved when . Smaller value can slow down the convergence of the algorithm, while bigger value can cause problems with over-correction. By this method, the input target of the GSW algorithm is modified according to the achieved target. And the non-uniformity caused by the imperfect optical system can be optimized.
3 Imaging and analyzing
The experimental setup is shown in Fig.2(a). A collimated 852nm laser beam illuminates the SLM (LCOS-SLM 15213-02, Hamamatsu Photonics) with a diameter of 10mm. Then the wavefront modulated beam passes through a Relay lens system (L1 and L2) and the phase pattern is transferred to the back pupil plane of the objective lens (Mitutoyo G Plan APO 50X). We insert a polarizing beam splitter(PBS) after L2 to pick up a small amount of light for intensity monitoring (CCD JHEM506GM). The most of power goes through an objective lens and then generates the tweezer array in vacuum chamber for trapping the atomic array. The backward fluorescence is picked up by a dichroic mirror and collected by an electron-multiplying CCD(EMCCD). The spot size is 1 m in the vacuum chamber and is 50m on CCD.
@@ 3 Imaging and analyzing
Since the CCD is an 8-bit deep camera, we can observe the saw-tooth shape from the 2D intensity distribution drawing in Fig.2 (c). The pixel size of the CCD we are using is 3.45 m 3.45 m. So, the focus will only take 15 pixels and it is impossible to neglect the saw-tooth shapes (detailed in Fig.2 (b)). To avoid the effect of saw-tooth shapes on the estimation of intensity equalization, we use the two-dimensional Gaussian fitting method to fit the tweezer intensity.
The fitting formula is:
@@
Here we use the general two-dimensional Gaussian fitting, which can give out the intensity, beam waist, and ellipticity at the same time. In Eq.4, is the amplitude of the Gaussian fitting function, representing the square root of intensity. And , are the center coordinates of the Gaussian function. D is the offset of the Gaussian function, which represents the intensity of the background. In Eq.5, , is the standard deviation of the Gaussian function in the x and y directions, which represents the width of the waist in the x and y directions respectively. is the rotation angle of the Gaussian fitting function which can reflect the aberration problem.
@@
We evaluate the time cost for producing a 3232 intensity equalized array. From inducing a target input to producing the first phase pattern, the GSW algorithm runs 20 cycles and takes less than 1 ms. Then the feedback process runs 10 cycles. Each feedback cycle contains several steps including finding out the position of the spots, 2D Gaussian fitting, intensity analysis, producing the new target, inducing the next GSW algorithm, and producing a new phase pattern. Each feedback process takes 23s. So, 4 minutes to 5 minutes is needed for producing an intensity equalized tweezer array with a size bigger than 1000.
4 Results
@@
Several algorithms can realize the intensity equalization. We test the uniformity of a 1010 tweezer array produced by different algorithms. According to the random mask(RM)[25], Gerchberg-Saxton(GS)[26], Generalized Adaptive Additive(GAA)[27] and GSW algorithms, the uniformity produced by GSW algorithm is the best. Table 1 is the test result of different strategies.
@@
The performance of different strategies is quantified by two parameters: uniformity (U) and percent standard deviation (). We use Eq.2. to character the uniformity of the array, where is the maximum intensity of the array while is the minimum intensity of array. Both are square of fitting results of A in Eq.4. And we use expression to character the RMS, where M represents the total number of optical holographic tweezers, represents the intensity of a single optical holographic tweezer, and represents the desired intensity of the optical holographic tweezer array.
Algorithm | Simulation | Experiment | ||
U() | () | U() | () | |
RM | 58.81 | 16.59 | 41.61 | 27.23 |
GS | 63.73 | 17.85 | 69.26 | 13.59 |
GAA | 77.86 | 9.12 | 73.56 | 9.31 |
GSW | 99.24 | 0.25 | 87.28 | 4.85 |
F-GSW | 98.96 | 0.47 |
@@
The experimental results show that for a 1010 array, the uniformity of simulated GSW is 99.24 and it is only 87.28 when directly loading the phase pattern on the SLM, which is due to the imperfect optical system. After our feedback optimization, the uniformity can be improved to 98.96. Compared to the theoretical limit of 99.24 calculated by the GSW algorithm, it is only 0.28 difference after feedback optimization, which means most of the non-uniformity caused by the optical system has been corrected.
@@
For the GSW algorithm, a random initial phase is given for iteration. This leads to the instability of the intensity equalization results. In the GSW algorithm we using, if the initial phase is fixed each time, then the phase diagram produced by the iteration is the same. Therefore, in the calculation process, we first use random initial phases for several GSW iterations and take the initial phase corresponding to the best uniformity as the initial phase of the feedback iteration. We call this the stationary phase method. We compare the uniformity results of the random initial phase and fixed initial phase, and the experimental results are shown in Fig.3. The uniformity stationary initial phase method is about 5 percentage points higher than the random phase method. And the error bar is much smaller. The error bars are calculated by 10 experiment runs.
@@
We evaluated the overall performance of the algorithm, and when the GSW algorithm was iterated 20 times, the algorithm results tended to be stable. When the camera feedback loop was done 10 times, the actual uniformity results tended to be stable. From the experimental results, we can find that the test uniformity of the GSW algorithm without image feedback tends to deteriorate when the array size increases. However, after the optimization of our feedback algorithm, the uniformity will be significantly improved.
5 Conclusion and outlook
@@
The neutral atomic array system is one of the most popular experimental systems in quantum computation and simulation. We built up this static tweezer system for trapping atomic arrays. To obtain an ideal single atomic array, the intensity equalization algorithm of the static tweezer array system is developed. We find that the fixed initial phase GSW algorithm can reduce the fluctuation of array uniformity and obtain better uniformity. Therefore, by several random phase GSW algorithms, the best initial phase pattern corresponding to the best homogenization result is selected. For the feedback GSW algorithm next, the selected initial phase pattern is used for better uniformity. The results show that this method can improve the uniformity by about 5%. On the other hand, the tweezers in the focal plane of the holographic tweezer array only occupy several pixels, so the intensity distribution obtained by the CCD camera must be a saw-tooth shape. These saw-tooth shapes will lead to inaccurate uniformity results. Therefore, we use two-dimensional Gaussian fitting to obtain the intensity distribution. At the same time, two-dimensional Gaussian fitting can also give out the information of spot size uniformity, and ellipticity of the tweezer trap that provides more judgment methods for further system optimization. The results show that the uniformity can exceed 96% when the tweezer size is bigger than 1000. The realization of a high uniformity tweezer array is of great significance to the many-body physics based on the single atom array.
@@
Funding National Key R&D Program of China under Grant No. 2020YFA0309400, NNSFC under Grants No. 12222409, and the Key Research and Development Program of Shanxi Province (Grant No. 202101150101025).
Acknowledgment H. S. acknowledges financial support from the Royal Society Newton International Fellowship Alumni follow-on funding (AL201024) of the UK.
Disclosures The authors declare no conflicts of interest.
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.

### v3

@@
Abstract
@@
High degree of adjustability enables the holographic tweezer array a versatile platform for creating an arbitrary geometrical atomic array. In holographic tweezer array experiments, an optical tweezer generated by a spatial light modulator (SLM) usually is used as a static tweezer array. However, the alternating current (AC) stark effect generally induces the intensity difference of traps in terms of different light shifts. So, intensity equalization is an essential prerequisite for preparing a many-body system with individually controlled atoms. Here, we report an intensity equalization algorithm. In particular, we observe the non-uniformity of the tweezer array is below 1.1 when the array size is larger than 1000. Our analysis shows that by optimizing the hardware performance of the optical system, this uniformity could be further improved. Our work offers the opportunities for large-scale quantum computation and simulation with reconfigurable atom arrays.
I Introduction
The holographic tweezer array formed by SLM is a technique that uses interference and diffraction principles to reconstruct the desired optical arrays [1, 2]. There are several protocols to produce the tweezer array, such as, by digital micro-mirror devices (DMD) [3], by Talbot effect [4], by micro-lens[5], by metasurface [6]. Compared to the SLM, the diffraction efficiencies of DMD and meta-surface are limited, while the arbitrary editing capability is impossible for Talbot and micro-lens array. In contrast, the phase pattern of SLM can be set on demand [7, 8], which makes it flexible for producing arbitrary arrays and correcting the aberration [9, 10, 11, 12]. Therefore, the SLM has become a high-efficiency equipment for generating optical tweezer arrays [13, 14].
Recent decade has witnessed great progresses in holographic tweezer arrays of neutral atoms. A key idea is outlined as follows: When the tweezer waist is small enough, the collisional-blocked effect leads to the fact that one atom at most can be trapped in the tweezer. Combined with the rearrangement technique, two- and three-dimensional defect-free atomic arrays have been realized [15, 16]. This reconfigurable platform opens a new avenue to large-scale quantum computation [17] and simulation [18, 19], and quantum metrology [20].
Concretely, the tweezer arrays made by SLM are used for static tweezer that plays the role of trapping the initial atomic array for sorting and holding the sorted atomic array [21]. In fact, the trap depth is proportional to the trap intensity, and an intensity equalized trap array is necessary [21] to prepare an ideal atomic array. Otherwise, an inhomogeneous trap depth in the tweezer array will result in the difference of the loss probability of the trapped atom as well as the different light shifts. Here we report a new scheme of the intensity equalization enabled by feedback weighted-Gerchberg-Saxton (GSW) algorithm [22, 23]. Through the imaging of the generated array, we create the corrected target and feed it back to the GSW algorithm. By leveraging the iterative feedback process, the non-uniformity of the tweezer array can be reduced significantly. The results show that the non-uniformity can be below 1.1% when the array size is over 1000.
II Principle of Feedback GSW
@@
The GSW algorithm [22] is the core of our approach. The specific process of the GSW algorithm and feedback is shown in Fig.1.The algorithm is based on Fourier optics theory and iteratively adjusts the phase distribution on a SLM to achieve the desired intensity distribution on the focal plane. The GSW algorithm optimizes the uniformity and quality of the holographic optical tweezer array by introducing weighted factors. We first generate a desired target diagram. The SLM is placed at the back focal plane (x-plane) of the lens. The reflected wavefront is shaped by a computer-generated hologram (CGH) [24] displayed on the SLM, and then forms the tweezer array at the focal plane (u-plane) of the lens.
@@
According to the imaging principle of Fourier optics, the pattern in the x-plane can be calculated to be . Due to the phase-only spatial light modulator, here we assign the amplitude to 1 and keep the phase component for fast Fourier-transform(FFT). A new target graph can be obtained by this FFT. So we could get the intensity distribution . To approach the desired target, a weighted factor is calculated by the expression below:
@@
After running many iterations inside the algorithm, the calculated target will be close to the desired target we input.
In practice, a grating phase needs to be added to shift the target away from the zero-th order. And due to the imperfect optical system, the achieved target is not uniform as the calculated result. To correct this non-uniformity caused by the optical system, we apply another feedback scheme. The basic principle is modifying the input target of GSW according to the non-uniformity target achieved by the GSW algorithm. This so-called feedback GSW algorithm begins with imaging the i-th GSW algorithm and analyzing the non-uniformity by 2D Gaussian fit of each spot. The non-uniformity (standard deviation) is calculated respectively by
where K represents the total number of optical holographic tweezers, is the i-th intensity of each spot, and is the average intensity of 2D Gaussian fit intensities in the i-th iteration. The (i+1)-th intensity is calculated as
where is the reconstructed target intensity for (i+1)-th iteration, is the target intensity of last loop. G is an adjustable gain factor. Here we apply a constant number as the gain factor. It turns out that the best performance is achieved when . Smaller value can slow down the convergence of the algorithm, while bigger value can cause problems with over-correction. By this method, the input target of the GSW algorithm is modified according to the achieved target. And the non-uniformity caused by the imperfect optical system can be optimized.
III Imaging and analyzing
The experimental setup is shown in Fig.2(a). A collimated 852 nm laser beam illuminates the SLM (LCOS-SLM 15213-02, Hamamatsu Photonics) with a diameter of 10 mm. Then the wavefront modulated beam passes through a Relay lens system (L1 and L2) and the phase pattern is transferred to the back pupil plane of the objective lens (Mitutoyo G Plan APO 50X). We insert a polarizing beam splitter (PBS) after L2 to pick up a small amount of light for intensity monitoring by a charge coupled device (CCD) camera (JHEM506GM). The most of power goes through an objective lens and then generates the tweezer array in vacuum chamber for trapping the atomic array. The backward fluorescence is picked up by a dichroic mirror and collected by an electron-multiplying CCD (EMCCD). The spot diameter is 1 m in the vacuum chamber and is 50 m on CCD.
@@ III Imaging and analyzing
Since the CCD is an 8-bit deep camera, we can observe the saw-tooth shape from the 2D intensity distribution drawing in Fig.2 (c). The pixel size of the CCD we are using is 3.45 m 3.45 m. So, the focus will only take 15 pixels and it is impossible to neglect the saw-tooth shapes (detailed in Fig.2 (b)). To avoid the effect of saw-tooth shapes on the estimation of intensity equalization, we use the two-dimensional Gaussian fit method to fit the tweezer intensity.
The fit formula is:
@@
Here we use the general two-dimensional Gaussian fit to determine the intensity, beam waist, and ellipticity simutanously. In Eq.4, is the amplitude of the Gaussian fit function, representing the square root of intensity. And , are the center coordinates of the Gaussian function. is the offset of the Gaussian function, which represents the intensity of the background. In Eq.5, and are the standard deviation of the Gaussian function in the x and y directions, which represents the width of the waist in the x and y directions respectively. is the rotation angle of the Gaussian fit function which can reflect the aberration problem.
@@
We evaluate the time cost for producing a 3232 intensity equalized array. From inducing a target input to producing the first phase pattern, the GSW algorithm runs 20 cycles and takes less than 1 ms. Then the feedback process runs 10 cycles. Each feedback cycle contains several steps including finding out the position of the spots, 2D Gaussian fit, intensity analysis, producing the new target, inducing the next GSW algorithm, and producing a new phase pattern. Each feedback process takes 23 s. Hence, a total time of 4 min to 5 min is needed for producing an intensity equalized tweezer array with a size larger than 1000.
IV Results
@@
Several algorithms can realize the intensity equalization. We test the non-uniformity of a 3232 tweezer array produced by different algorithms. According to the Gerchberg-Saxton (GS) [25], Generalized Adaptive Additive (GAA) [26] and GSW algorithms, the non-uniformity produced by the GSW algorithm is the least. Table 1 is the test result of different strategies.
@@
The performance of different strategies is quantified by non-uniformity (standard deviation,). We use Eq.2. to character the non-uniformity of the array, where K represents the total number of optical holographic tweezers, is the i-th intensity of each spot.
| Algorithm | Simulation | Experiment |
| GS | 18.18 | 24.84 |
| GAA | 14.81 | 24.69 |
| GSW | 0.44 | 11.007 |
| F-GSW | - | 1.006 |
@@
The experimental results show that for a 3232 array, the non-uniformity of simulated GSW is 0.44 and it is 11.007 when directly loading the phase pattern on the SLM, which is due to the imperfect optical system. After our feedback optimization, the non-uniformity can be reduced to 1.006. Compared to the theoretical limit of 0.44 calculated by the GSW algorithm, it is only 0.566 difference after our feedback optimization, which means most of the non-uniformity caused by the optical system has been corrected.
@@
For the GSW algorithm, a random initial phase is given for iteration. This leads to the instability of the intensity equalization results. In the GSW algorithm we using, if the initial phase is fixed each time, then the phase diagram produced by the iteration is the same. Therefore, in the calculation process, we first use random initial phases for several GSW iterations and take the initial phase corresponding to the best uniformity as the initial phase of the feedback iteration. We call this the stationary phase method. The purpose of this process is to select the most effective phase to improve the uniformity of the holographic tweezers array rather than relying on more iterations to optimize the experimental results.We compare the non-uniformity results of the random initial phase and fixed initial phase, and the experimental results are shown in Fig.3. The non-uniformity of stationary initial phase method is about 1.5 percentage points lower than that of the random phase method. And the error bar is much smaller. The error bars are calculated by 10 experiment runs.
@@
We evaluated the overall performance of the algorithm, and when the GSW algorithm was iterated 20 times, the algorithm results tended to be stable. As shown in Fig.4, the initial standard deviation increases as the array size increases.And when the camera feedback loop was done 10 times, the actual uniformity results tended to be stable. From the experimental results, we can find that the test uniformity of the GSW algorithm without image feedback tends to deteriorate when the array size increases.To present the experimental results of intensity equalization, we analyzed the final array results of the 32×32 array in Fig.4. Fig.5 (a) is a graph of normalized intensity data, which is the date acquired from Thorlabs’ CMOS Camera Beam Profiler (BC106N - VIS/M). Fig.5(b) shows the statistical results of the tweezer array intensities. It can be seen from the results that the uniformity of the array has been significantly improved by the feedback intensity homogenization algorithm.
V Conclusion and outlook
@@
The neutral atomic array system is one of the most popular experimental systems in quantum computation and simulation. We built up this static tweezer system for trapping atomic arrays. To obtain an ideal single atomic array, the intensity equalization algorithm of the static tweezer array system is developed. We find that the fixed initial phase GSW algorithm can reduce the fluctuation of array uniformity and obtain better uniformity. Therefore, by several random phase GSW algorithms, the best initial phase pattern corresponding to the best homogenization result is selected. For the feedback GSW algorithm next, the selected initial phase pattern is used for better uniformity. The results show that this method can reduce the non-uniformity by about 1.1%. On the other hand, the tweezers in the focal plane of the holographic tweezer array only occupy several pixels, so the intensity distribution obtained by the CCD camera must be a saw-tooth shape. These saw-tooth shapes will lead to inaccurate uniformity results. Therefore, we use two-dimensional Gaussian fit to obtain the intensity distribution. At the same time, two-dimensional Gaussian fit can also give out the information of spot size uniformity, and ellipticity of the tweezer trap that provides more judgment methods for further system optimization. The results show that the non-uniformity can below 1.1% when the tweezer size is bigger than 1000. The realization of a high uniformity tweezer array is of great significance to the many-body physics based on the single atom array.
Acknowledgements
@@
This work was supported by National Key R&D Program of China under Grant No. 2020YFA0309400, NNSFC under Grant No. 12222409, and the Key Research and Development Program of Shanxi Province (Grant No. 202101150101025).